OpenStudy (anonymous):

I think I have the right answer, can you please check? Question is in comments.

OpenStudy (anonymous):

\[\frac{3x^4}{y^3}\div \frac{6x^2}{y^2}\]

OpenStudy (anonymous):

thats the question, heres what I think the answer is.

OpenStudy (anonymous):

\[\frac{ 2x^2 }{ y }\]

OpenStudy (anonymous):

Say to simplify by division.

OpenStudy (primeralph):

No, small error.

OpenStudy (anonymous):

what is it?

OpenStudy (primeralph):

try again...............

OpenStudy (anonymous):

ok, lets see.

OpenStudy (anonymous):

but we're not multiplying.

OpenStudy (luigi0210):

You canceled wrong

OpenStudy (primeralph):

turn the one on the right upside down and simply multiply

OpenStudy (anonymous):

\[\frac{ x^2 }{ 2y }\] Is the only other plausible option.

OpenStudy (primeralph):

yeah

OpenStudy (anonymous):

The other option has 18x^2 as the numerator, and thats multiplying.

OpenStudy (anonymous):

ok, thanks, I have another one, one second...

OpenStudy (primeralph):

Make a new post. and gimme my medal!! lol

OpenStudy (luigi0210):

Gottcha